OpenStudy (macabbey101):

PLEASE HELP WILL FAN AND MEDAL!!!!!!!!! Write the point-slope form of the equation of the line through the given point with the given slope. (-4,3) slope= -1 A) y+3=1/5(x+4) B) y-3=-(x+4) C) y-3=-2(x+4) D) y+3=3(x-4)

OpenStudy (nikkibxox):

You want to find the equation for a line that passes through the point (-4,3) and has a slope of -1. First of all, remember what the equation of a line is: y = mx+b Where: m is the slope, and b is the y-intercept To start, you know what m is; it's just the slope, which you said was -1. So you can right away fill in the equation for a line somewhat to read: y=-1x+b. Now, what about b, the y-intercept? To find b, think about what your (x,y) point means: (-4,3). When x of the line is -4, y of the line must be 3. Because you said the line passes through this point, right? Now, look at our line's equation so far: . b is what we want, the -1 is already set and x and y are just two "free variables" sitting there. We can plug anything we want in for x and y here, but we want the equation for the line that specfically passes through the the point (-4,3). So, why not plug in for x the number -4 and for y the number 3? This will allow us to solve for b for the particular line that passes through the point you gave!. (-4,3). y=mx+b or 3=-1 × -4+b, or solving for b: b=3-(-1)(-4). b=-1. The equation of the line that passes through the point (-4,3) with a slope of -1 is y=-1x-1 Is that all of the question? http://www.webmath.com/equline2.html

OpenStudy (macabbey101):

oh my gosh thank you so much! I understand it now. you helped me so much!!!